OpenStudy (anonymous):

Super lost on homework question!! Distribute, simplify, multiply by 100 and solve for x graph. π(x³ + 4.5x²)-14.13=0

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

\[\pi x^3+4.5\pi x^2-14.13=0\] \[100\pi x^3+450\pi x^2-1413=0\] after that i am stuck

OpenStudy (anonymous):

algebra won't do it for you http://www.wolframalpha.com/input/?i=100 \pi+x^3%2B450\pi+x^2-1413%3D0

OpenStudy (anonymous):

thanks for trying. I really don't know what I'm supposed to do, this is for polynomial functions homework. I've already tried that site, I'm pretty sure it doesn't give the correct answer.